

********************************************************************
UARS WEEKLY STATUS REPORT - 08 Nov 2002
********************************************************************
***********
SOLSTICE
***********
During this week, SOLSTICE continued to operate in good health gathering
solar data on all available orbital tracking periods.
***********
SUSIM
***********
SUSIM has successfully performed its daily mid-resolution and MgII
solar scans this past week during periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Notheing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is currently operation. It required a re-set today, after an
error report, but otherwise is operating normally.
***********
MPG
***********
The MPG assisted in troubleshooting the problem
of DECNET communications between instrument Remote
Analysis Computers and the UARS Command Management
System.
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 October 2002 (Orbit 60835, GMT Day 299) through
01 November 2002 (Orbit 60939, GMT Day 305).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 301/2057 60878 POWER ON
PEM 301/2057 60878 POWER ON (DAY-ONLY)
SUSIM 301/2242 60879 POWER ON
WINDII 302/1304 60888 POWER ON
PEM 305/1300 60933 POWER ON AXIS 1 & 2
The beta angle ranged between the angles of 64.3 degrees to 68.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 19 November 2002 (DOY 323).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.25 to 1.26 degrees C. Additional
battery performance data for each battery on 25 October and 01
November 2002 are provided in Appendix D.
The clock error ranged from -10.4 to 14.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 13
PEM = 02
SOLSTICE = 01
SUSIM = 01
WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT is happy to report that we have successfully recovered from
the OBC halt anomaly which occurred on 21 October 2002. All
instruments that are scheduled on are now back on with the exception
of HRDI. The FOT observed a HRDI HPC protection telemetry point
toggling at a fairly rapid pace before the anomaly. The HRDI PI team
was notified of this and investigated this circuitry. They found that
this signature occurs nominally and do not believe this to have caused
the OBC halt. Data is being pulled from earlier time frames to verify
this and if this signature is found, HRDI will be turned back on first
thing next week.
A full analysis of the OBC hardware dumps, which were taken
immediately after the anomaly, has not revealed any indications of any
abnormal behavior. The FOT is planning to try to load this image to
the simulator in order to discover more about the state of the OBC
during the anomaly. If this test does not show any clues about the
problem then the cause of the anomaly may remain a question mark (or
be attributed to a single event upset). The investigation continues
but as of right now there are no firm leads to follow.
MEETINGS
The FOT attended several meetings throughout the week to discuss
the safehold recovery activities.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 November 2002 (DOY 322).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 November 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60840 299/0944 02-405 OBC GYRO BIAS RESET
60854 300/0718 02-405 OBC GYRO BIAS RESET
60878 301/2057 02-411 SOLSTICE POWER ON
60878 301/2057 02-412 PEM POWER ON (DAY-ONLY)
60878 301/2138 02-405 OBC BIAS ADJUST
60879 301/2242 02-413 SUSIM POWER ON (BEGIN)
60879 301/2316 02-413 SUSIM POWER ON (COMPLETE)
60888 302/1304 02-414 WINDII POWER ON
60891 302/1914 02-415 OBC/CLK CLOCK FIX
60893 302/2106 02-416 MACS ESA-2 OFF/HTR ON
60894 302/2241 02-417 OBC TABLE DUMPS (23 & 66)
60897 303/0353 02-418 OBC/CLK CLK ADJ 0.0 TO 20.6 MSEC/DAY
60899 303/0755 02-419 OBC/CLK CLK ADJ 20.6 TO 41.2 MSEC/DAY
60904 303/1549 02-420 SSPP/HGA SSPP & HGA TO ACTUAL MODE /RWM ENA
60907 303/1954 02-421 MACS FOUR WHEEL CONTROL ENABLED
60909 303/2240 02-405 OBC GYRO BIAS ADJUST
60933 305/1300 02-422 PEM POWER ON AXIS 1 & 2
60936 305/1810 02-423 OBC CLK ADJ FR 41.2 TO 20.6 SEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 November 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 November 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60897 303/0358 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
60899 303/0819 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
60936 305/1823 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 November 2002
25 October 2002 - GMT Day 298 Beta = 62.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -145.6/-168.0 -2.42/-3.68 +00.0/+00.0 08.3 0.000 0.0
2 +00.0/-33.6 +3.55/+2.60 +20.8/-14.4 27.0 1.050 14.9
3 +22.4/-28.0 +2.60/+1.65 +20.0/-14.4 27.0 1.040 14.7
01 November 2002 - GMT Day 304 Beta = 68.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -145.6/-179.2 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +05.6/-33.6 +3.23/+1.97 +19.2/-16.0 26.6 1.040 15.5
3 +16.8/-22.4 +2.60/+1.34 +18.4/-15.2 26.4 1.030 15.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
- None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 November 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-052 SOLSTICE 60366 267/1951 N/A LOADING PROBLEM
02-054 FOT 60689 289/0819 31607 FORWARD FAIL-OVER AT WHITESAND
20 SEC RTD lost, non-recover
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-053 FOT 60630 285/1008 30382 GENERIC LATE ACQUISITION
02-056 FOT 60775 295/0122 OBC CRASH